myininaya (myininaya):

ok who can integrate f'/(f-f') with respect to x and yes f is a function of x i can't think right now

OpenStudy (anonymous):

is that related to other problem (f g)'=(f'g') ?

myininaya (myininaya):

yes

myininaya (myininaya):

i don't think we can do it by any elementary ways

OpenStudy (anonymous):

did you try \[\int \frac{x^3}{x^3-3x^2}dx\] or \[\int\frac{\sin(x)}{\sin(x)-\cos(x)}dx\]?maybe there is a pattern?

myininaya (myininaya):

maybe i should look for a pattern i will be back

OpenStudy (anonymous):

well i tried them both with wolfram. then i went to http://www.wolframalpha.com/input/?i=intgerate+f%28x%29%2F%28f%28x%29-f%27%28x%29%29dx

OpenStudy (anonymous):

results not promising

myininaya (myininaya):

sorry the top should be f' too ok for real though i will be right back

OpenStudy (anonymous):

now you tell us !

myininaya (myininaya):

lol no it was f' all the long satellite you are getting old

myininaya (myininaya):

i see a pattern with polynomials or monomials anyways

myininaya (myininaya):

no not a pattern

myininaya (myininaya):

just one for monomials

myininaya (myininaya):

i give up i need to do my own work now

OpenStudy (anonymous):

A simple answer occurs to me, \[\int\limits_{}^{}f \prime/(f-f \prime) = \ln(f-f \prime ) \] if and only if \[f \prime \prime=0\] so formula will only work for monomials like you just said

OpenStudy (anonymous):

wolfram says its a nonlinear differential equation, had the same feeling when I saw it, http://www.wolframalpha.com/input/?i=integrate+a%27%2F%28a-a%27%29